the total area of a sphere is 144 pie square centimeters. what is the diameter?
Answer by nerdybill(7384)

the total area of a sphere is 144 pie square centimeters. what is the diameter?
.
since the area of a sphere is given as:
4(pi)r^2
.
We set the above equal to 144 to find the radius:
4(pi)r^2 = 144(pi)
dividing both sides by pi:
4r^2 = 144
dividing both sides by 4:
r^2 = 36
take the square root of both sides:
r = 6 cm
.
But, the problem wanted the diameter...
diameter = 2r
diameter = 2(6)
diameter = 12 cm (answer)